# Eating Speed Measurement Using Wrist-Worn IMU Sensors in Free-Living Environments

### v1

@@ 5.1.1 Bite Detection
The bite detection in free-living settings was tested on FD-I dataset. The 7-fold cross validation was used on FD-I dataset. The split of this dataset was on participant-level to avoid information leakage between the train set and test set. As mentioned in Section III-C, the FD-I dataset is highly unbalanced, hence, the MO and OREBA datasets were included as part of the training set. Results are shown in Table II. For point-wise results, TCN-MHA obtained the highest kappa score (0.735). For segment-wise results, the TCN-MHA also achieved the highest F1-score for eating with the value of 0.849 and 0.781, for $k=0.1$ and $k=0.5$, respectively. However, MS-TCN yielded a higher F1-score for drinking compared to TCN-MHA (0.906$\rightarrow$0.902, for $k=0.1$).
@@ 5.1.2 Eating Episode Detection
The predicted bite sequences from previous step were clustered into eating episodes using the DBSCAN-based algorithm. The results of eating episodes detection are shown in Table III. Among the 74 ground truth eating episodes, the TCN-MHA successfully detected 64 sessions with a mean IoU of 0.899.
@@ 5.1.3 Eating Speed
For eating speed estimation, the MAPE and PCC are shown in Table III. The TCN-MHA model had the least MAPE of 0.110; the highest PCC value of 0.925. The scatter plots are drawn to graphically demonstrate the correlation between predicted and ground truth eating speed in the FD-I dataset (Fig. 7). Most eating episodes had an eating speed falling into the range of 2-6 bite/min. Additionally, the result for each individual episode on TCN-MHA model has been shown in Fig. 8.
Figure 9: The scatter plot for eating speed on FD-II dataset.
5.2 Experiments on FD-II Dataset
To further validate the proposed method on eating speed measurement, the FD-II was used as the hold-out dataset. We utilized two in-meal datasets (OREBA, MO) and the FD-I dataset to train our model, then used the entire FD-II as the test set. It should be noted that we were only able to measure the eating speed during meal sessions, as we lack labels for out-of-meal sessions. Hence, we did not evaluate the predicted snack sessions (eating episodes with duration less than 7 min). The performance of in-meal bite detection, meal detection, and speed measurement are shown in Table IV. For in-meal bite detection, the TCN-MHA model yielded the highest segmental F1-score with 0.820 when $k=0.1$. The TCN-MHA obtained the best performance in meal detection, which successfully detected 48 meal sessions (7 FPs, 4 FNs), with an F1-score of 0.897 and a mean IoU of 0.841. For eating speed measurement, the TCN-MHA had the MAPE of 0.146 and the PCC of 0.924. The ResNet-BiLSTM had the highest number of TP meals, however the number of FPs was also higher than that of the TCN-MHA. Scatter plots to show the deviation between predicted and true eating speed (only for meal sessions) in the FD-II dataset are shown in Fig. 9.
TABLE V: Existing studies on Food Intake Monitoring in Free-living Environment
[…]
@@ 6 Discussion
In this study, we first tested several models on free-living datasets (FD-I and FD-II) for bite detection. Results from Table II show that bite detection in free-living environments is feasible. Meanwhile, Table III also indicates a tendency for models to overlook certain eating episodes, as evidenced by a higher number of FNs compared to FPs. Upon comparing the output and the annotation video, we found that FNs are mainly from snack sessions, implying that the clustering of snack sessions is more challenging than detecting meals. One reason is that the eating patterns of snacking can vary widely in terms of frequency and duration, making it difficult to cluster all snacking gestures into specific episodes. Other eating episode detection studies also suffer from this [10, 38].
@@
Figure 10: The 2 h IMU data segment from both hands and the corresponding bite and meal detection examples. Additionally, a 2-min in-meal segment is also selected to show the detailed bite detection.
Figure 11: The minute-level eating speed distribution from one participant in FD-I dataset.
@@
When comparing our work to prior studies on food intake monitoring in free-living environments (as shown in Table V), the dataset exhibits a comparable size to others, except for the dataset used by Sharma et al. [10, 44], which is substantially larger than the rest. However, it’s worth noting that existing free-living datasets mainly focus on eating episode detection, which only requires labeling the starting and ending times of eating episodes. In contrast, our data for eating speed measurement requires bite-level annotation, which demands additional efforts for data collection and annotation. The detection granularity of this study is shown in Fig. 10. Meanwhile, compared to eyeglass-based and necklace based approaches focusing on mouth-throat movements for food intake monitoring, one limitation of wristband-based approach is that we need to wear IMUs on both hands. Wearing the IMU only on the dominant hand may lead to the omission of some eating or drinking gestures in free-living environments, as illustrated in Fig. 2. However, the advantage of the wrist-worn IMU sensor lies in its wide integration into smartwatch products, making it more readily accepted by users compared to other solutions. Table V can also show the prevalence of wrist-worn IMU.
In food intake monitoring, several wrist-worn IMU based bite detection datasets have been published, however, they are only used separately to benchmark performance of different models. In this study, we examined the feasibility of combining different datasets for training. When performing bite detection on the FD-I dataset, our own MO dataset and the external public OREBA dataset were included as part of training set. This is the first attempt to integrate different datasets for food intake monitoring, resulting in a 0.7% increase of F1-score in bite detection, 3.4% increase of PCC in eating speed detection. However, it is important to note that the prerequisite of this integration is that the coordinates of the IMU sensors used in both datasets are the same. The coordinates of IMU in OREBA datasets were adjusted to make sure the orientation of $x$, $y$, $z$ axis of the two datasets were consistent (ours is depicted in Fig. 1). Meanwhile, the two-hand combination method also facilitates the training of multiple datasets, as it allows one-hand datasets and two-hands datasets. Specifically, to handle data from both hands in free-living environments, we applied the hand mirroring + temporal concatenation method to combine IMU data from two hands. This solution also remains flexibility to process data from single hand IMU.
@@
To assess the model’s complexity, the number of parameters for each model and their floating point operations (FLOPs) are indicated in Table VI. The number of parameters in ResNet-(Bi)LSTM-based models were significantly larger than CNN-LSTM (0.134 M) and TCN-MHA models (0.203 M). Additionally, a test was carried out to assess the latency for processing 1 min data using the TCN-MHA model. Utilizing a laptop equipped with an Intel Core i7 10750 CPU @2.6 GB, 6 cores (no GPU configuration), the TCN-MHA required 37.61 ms to generate predictions.
TABLE VI: Model Complexity
Model #Paras (M) # FLOPs (G)
CNN-LSTM 0.134 0.129
CNN-BiLSTM 0.241 0.233
ResNet-LSTM 3.078 2.955
ResNet-BiLSTM 3.415 3.280
MS-TCN 0.298 0.284
TCN-MHA 0.203 0.194
TABLE VII: Minute-level Eating Speed
@@
The proposed wrist-worn IMU-based approach has the potential to replace the questionnaire-based eating speed surveys in nutrition studies, which offers more accurate quantitative results, thereby advancing the analysis of the correlation between eating speed and obesity-related problems. Additionally, such an eating behavior-related digital biomarker can be applied to individuals interested in documenting their daily dietary habits for long-term dietary health assessment.
@@
This study mainly focus on eating speed measurement for each eating episode, resulting in an averaged speed per episode, which is a well-accepted definition. Additionally, we also explored the feasibility of measuring minute-level eating speed. To achieve this, the number of detected bites in each minute is considered as minute-level eating speed. Fig. 11 shows the minute-level eating speed distribution through one day. Results are shown in Table VII. The TCN-MHA had the MAPE of 0.181 and the PCC of 0.840 on FD-I, and had the MAPE of 0.212 and the PCC of 0.834 on FD-II. The meal-level eating speed represents a holistic view of an eating session, whereas minute-level speed provides insights into eating patterns at a more granular level and is suitable for studying immediate speed changes during a meal. On the other hand, the performance of minute-level speed detection (Table VII) implies that the minute-level eating speed detection is more challenging compared to meal-level speed detection.
@@
This research utilized IMU wristband to measure eating speed in free-living environment. The results were promising, however, it should be noted that the proposed method were based on off-line processing. To maximize its application scenarios, it is worthwhile to exploit the feasibility of implementing this method to smartwatch-smartphone setup for daily eating speed monitoring. Another limitation is that obtaining fine-annotated ground truth data is troublesome. In our approach, research assistants had to follow participants activities to record videos. Existing wearable-based cameras [13] can be used for recording to minimize the effort in future study. Furthermore, to quantify the actual food intake, the wearable IMU system has the potential to be combined with the smart plate [16] and the smart snack box [45] to estimate the calorie intake in real life.
@@ 7 Conclusion
In this work, we presented a comprehensive framework for automated measurement of eating speed in free-living environments. To the best of our knowledge, this is the first of its kind. This framework has the potential to extend the application scope of the automated food intake monitoring field. The framework mainly relies on two essential parts: bite detection and eating episode detection in free-living environments. The success of bite detection paves the way for eating episode detection and segmentation, resulting in a good capability for eating speed measurement. The hold-out experiments further underscore its robustness in meal-level eating speed detection.

### v2

@@ 5.1.1 Bite Detection
The 7-fold cross validation was used on FD-I dataset. Specifically, within each fold, all corresponding IMU data from participants designated in the test list were allocated to the test set to avoid information leakage. As mentioned in Section III-C, the FD-I dataset is highly unbalanced, hence, the MO and OREBA datasets were included as part of the training set. Results are shown in Table 2. For point-wise results, TCN-MHA obtained the highest kappa score (0.735). For segment-wise results, the TCN-MHA also achieved the highest F1-score for eating with the value of 0.849 and 0.781, for $k=0.1$ and $k=0.5$, respectively. However, MS-TCN yielded a higher F1-score for drinking compared to TCN-MHA (0.906$\rightarrow$0.902, for $k=0.1$).
@@ 5.1.2 Eating Episode Detection
The predicted bite sequences from previous step were clustered into eating episodes using the DBSCAN-based algorithm. The results of eating episodes detection are shown in Table 3. Among the 74 ground truth eating episodes, the TCN-MHA successfully detected 64 sessions with a mean IoU of 0.899.
@@ 5.1.3 Eating Speed
Table 3 shows the eating speed estimation results. The TCN-MHA model had the least MAPE of 0.110; the highest PCC value of 0.925. The scatter plots are drawn to visualize the correlation between predicted and ground truth eating speed in the FD-I dataset (Fig. 10). Most eating episodes had an eating speed falling into the range of 2-6 bite/min. Additionally, the result for each individual episode on TCN-MHA model has been shown in Fig. 11.
Table 7: Existing studies on Automated Food Intake Monitoring in full-day scenarios
Work Positiona Sensorb # Participants # Days # Hours Eating Episode Bite annotation Eating
detection & detectionc speed
@@
5.2 Experiments on FD-II Dataset
To further validate the proposed method on eating speed measurement, the FD-II was used as the holdout dataset. We utilized two in-meal datasets (OREBA, MO) and the FD-I dataset to train our model, then used the entire FD-II as the test set. It should be noted that we were only able to measure the eating speed during meal sessions, as we lack labels for out-of-meal sessions. Hence, we did not evaluate the predicted snack sessions (eating episodes with duration less than 7 min). The performance of in-meal bite detection, meal detection, and speed measurement are shown in Table 4. For in-meal bite detection, the TCN-MHA model yielded the highest segmental F1-score with 0.820 when $k=0.1$. The TCN-MHA obtained the best performance in meal detection, which successfully detected 48 meal sessions (7 FPs, 4 FNs), with an F1-score of 0.897 and a mean IoU of 0.841. For eating speed, the TCN-MHA had the MAPE of 0.146 and PCC of 0.924. The ResNet-BiLSTM had the highest number of TP meals, however the number of FPs was also higher than that of TCN-MHA.
5.3 Experiments on Public Datasets
To compare the performance across datasets, the public OREBA and Clemson datasets were used as benchmarks. Both 7-fold and holdout validation methods were employed. In the 7-fold validation, the split of test set in each fold was participant-level to avoid information leakage. The TCN-MHA model was trained and evaluated within each dataset separately. For holdout validation, the model was trained on FD-I and MO datasets and then tested on OREBA and Clemson, respectively. Results are shown in Table 5. In 7-fold validation, it is evident that higher F1-scores for eating gesture detection and episodes detection were obtained on the OREBA and Clemson datasets compared to that of FD-I dataset (Table 2 and Table 3). The OREBA achieves the lowest MAPE of 0.069 with the PCC of all datasets exceeding 0.900. These results validate the performance of proposed approach, implying that it is indeed more challenging to detect eating gestures and episodes in near-free-living environments. Results also show that measuring eating speed in an uncontrolled environment (i.e., cafeteria) is more challenging than in a controlled environment (i.e., lab). In holdout validation, all evaluation metrics decreased compared to those in the 7-fold validation. On OREBA dataset, the MAPE of eating speed increased from 0.069 to 0.119. The reduction on the Clemson was much larger than that on OREBA dataset, the MAPE on Clemson increased from 0.097 to 0.264.
Table 8: Dominant hand performance on FD-I dataset using TCN-MHA
Hands Eating Gesture F1-score Eating Episodes Detection Eating Speed
$k=0.1$ $k=0.5$ TP FP FN F1 IoU MAPE PCC
Both hands 0.849 0.781 64 0 10 0.928 0.899 0.110 0.925
Dominant hand 0.806 0.738 56 0 18 0.862 0.857 0.128 0.854
@@ 6 Discussion
In this study, we first tested several models on FD-I and FD-II for bite detection. Results from Table 2 show that bite detection in near-free-living environments is feasible. Meanwhile, Table 3 also indicates a tendency for models to overlook certain eating episodes, as evidenced by a higher number of FNs compared to FPs. Upon comparing the output and the annotation video, we found that FNs are mainly from snack sessions, implying that the clustering of snack sessions is more challenging than detecting meals [10, 40]. One reason is that the eating patterns of snacking can vary widely in terms of frequency and duration, making it difficult to cluster all snacking gestures into specific episodes.
A deeper analysis has also been conducted to compare the model’s performance across different eating styles and environments, as shown in Table 6. Regarding four eating styles, it is discerned that eating with hand exhibits the lowest performance in terms of both Recall and IoU scores in episodes detection, as well as the MAPE in eating speed measurement step. This implies that the model tends to miss hand eating episodes. Based on observation, we found that eating with hand is more difficult to detect. This difficulty is attributed to the inherent variability associated with this eating style. Additionally, hand-eating gestures closely resemble other hand movements commonly encountered in daily life, further complicating the detection. The challenge of detecting snacking sessions is also linked to such issue. Among three types of eating locations, eating in restaurants has shown the highest performance by detecting all episodes and achieving the highest IoU score. One potential reason is that individuals tend to eat in a more formalized manner in restaurant, often using utensils such as fork & knife. By checking the data, the FNs in workspace are mainly snacking sessions.
@@
Figure 12: The 100 min IMU data segment from both hands and the corresponding bite and meal detection examples. Additionally, a 2-min in-meal segment is also selected to show the detailed bite detection.
@@
When comparing our work to prior studies in full-day scenarios (as shown in Table 7), the dataset exhibits a comparable size to others, except for the dataset used by Sharma et al. [10, 46], which is substantially larger than the rest. However, it’s worth noting that existing free-living datasets mainly focus on eating episode detection, which only requires labeling the starting and ending times of eating episodes. In contrast, our data for eating speed measurement requires bite-level annotation, which demands additional efforts for data collection and annotation. The detection granularity of this study is shown in Fig. 12. Compared to eyeglass-based and necklace based approaches, one limitation of wristband-based approach is that we need to wear IMUs on both hands. Wearing the IMU only on the dominant hand may lead to the omission of some eating gestures, as illustrated in Table 8. In comparison to utilizing IMU data from both hands, #FNs increased from 10 to 18 due to the absence of bites from non-dominant hand, the MAPE increased from 0.110 to 0.128. These results suggest that using IMU data from only the dominant hand leads to missed detections of eating episodes. Additionally, for the eating speed from detected episodes, the MAPE performance also reduced. However, considering that people typically wear only one watch in normal daily life, the performance may still be acceptable.
Several wrist-worn IMU based bite detection datasets have been published, however, they are only used separately to benchmark performance of different models. In this study, we examined the feasibility of combining different datasets for training. When performing bite detection on the FD-I dataset, our own MO dataset and the external public OREBA dataset were included as part of training set. This is the first attempt to integrate different datasets for food intake monitoring, resulting in a 0.7% increase of F1-score in bite detection, 3.4% increase of PCC in eating speed detection. The 7-fold validation results on Clemson and OREBA datasets showed in Table 5 demonstrate that the proposed approach performs better within meal-session datasets. Whereas the holdout validation results exhibit lower performance, especially for the Clemson dataset. One potential reason could be attributed to the orientation of the IMU sensor used in Clemson, which differs from that of the FD-I and OREBA. This suggests that the generalization ability of proposed approach may be limited by the orientation of IMU. Table 5 also shows that the largest dataset (Clemson) exhibits the highest variability among meal-session datasets, in terms of number of subjects, types of foods, and types of behaviors. This variability is also illustrated in the ground truth plots of eating speed in Fig. 4.
Regarding the scatter plot in Fig. 10, we noticed a common outlier point in CNN-(Bi)LSTM and ResNet-(Bi)LSTM models, representing an eating episode with a ground truth speed of 8.54 bites/min. For ResNet-BiLSTM, the predicted speed is 2.98 bites/min. By inspecting the video, we found this outlier episode involved eating snacks with hand. This observation aligns with findings from eating style analysis, suggesting that hand-eating gestures are more challenging to detect.
Figure 13: The minute-level eating speed distribution in a full day from one participant in FD-I dataset.
@@
To assess the model’s complexity, the number of parameters for each model and their floating point operations (FLOPs) are indicated in Table 4. The number of parameters in ResNet-(Bi)LSTM-based models were significantly larger than CNN-LSTM (0.134 M) and TCN-MHA models (0.203 M). Additionally, a test was carried out to assess the latency for processing 1 min data using the TCN-MHA model. Utilizing a laptop equipped with an Intel Core i7 10750 CPU @2.6 GB, 6 cores (no GPU configuration), the TCN-MHA required 37.61 ms to generate predictions.
Table 9: Minute-level Eating Speed
@@
The proposed wrist-worn IMU-based approach has the potential to replace the questionnaire-based surveys on eating speed in nutrition studies, advancing the analysis of the correlation between eating speed and obesity-related problems. Moreover, such a dietary-related digital biomarker can be applied to individuals interested in documenting their daily dietary habits. Further, combining data from fitness trackers, dietary analysis apps, and physiological sensors can generate a more comprehensive picture of an individual’s health status, facilitating a holistic understanding of individual health profiles. For diabetes management, such technology can be integrated with the wearable continuous glucose monitor (CGM) sensors to gain deeper insights into postprandial glucose fluctuations and their correlation with eating behaviors. By investigating correlations between eating speed, meal composition, and postprandial glucose levels, predictive algorithms could be developed to anticipate glycemic responses to specific meals, thus empowering individuals with diabetes to proactively manage their blood glucose levels and mitigate the risk of hyperglycemia or hypoglycemia.
@@
This study mainly focus on eating speed measurement for each eating episode, resulting in an averaged speed per episode, which is a well-accepted definition. Additionally, we also explored the feasibility of measuring minute-level eating speed. To achieve this, the number of detected bites in each minute is considered as minute-level eating speed. Fig. 13 shows the minute-level eating speed distribution through one day. Results are shown in Table 9. The TCN-MHA had the MAPE of 0.181 and the PCC of 0.840 on FD-I, and had the MAPE of 0.212 and the PCC of 0.834 on FD-II. The meal-level eating speed represents a holistic view of an eating session, whereas minute-level speed provides insights into eating patterns at a more granular level and is suitable for studying immediate speed changes during a meal. On the other hand, the performance of minute-level speed detection (Table 9) implies that the minute-level eating speed detection is more challenging compared to meal-level speed detection.
@@
The collected dataset served to validate the feasibility of eating speed measurement in near-free-living environments. Participants were recorded in their habitual environments where they usually eat, work, and study, thereby ensuring a sense of familiarity with their surroundings. Nevertheless, one limitation is that the participants behaviors may potentially be affected when they were followed and recorded all day, which is the reason that we use the term near-free-living environments. Additionally, the dataset only contains conventional eating behaviors, namely structured meal consumption and inter-meal snacking. Other types of eating behavior such as grazing, eating while walking, and night eating, were not included in the study. A larger dataset with a broader spectrum of eating behaviors would be helpful to further validate the robustness of the approach in the future. Meanwhile, the proposed method were based on off-line processing. It is worthwhile to exploit the feasibility of implementing this method to smartwatch-smartphone setup for daily eating speed monitoring. Another limitation is that obtaining fine-annotated ground truth data is troublesome. In our approach, research assistants had to follow participants activities to record videos. Existing wearable-based cameras [13] can be used for recording to minimize the effort in future study. Furthermore, to quantify the actual food intake, the wearable IMU system has the potential to be combined with the smart plate [16] and the smart snack box [47] to estimate the calorie intake in real life.
@@ 7 Conclusion
In this work, we presented a comprehensive framework for automated measurement of eating speed in near-free-living environments. To the best of our knowledge, this is the first of its kind. This framework has the potential to extend the application scope of the automated food intake monitoring field. The proposed method is built upon two essential tasks in automated food intake monitoring domain: bite detection and eating episode detection. The success of bite detection paves the way for eating episode detection and segmentation, resulting in a good capability for eating speed measurement. The holdout experiments further underscore its robustness in meal-level eating speed detection.